A solar energy collector has dimension of 1.2m by 2.4m and collects solar energy with about 70% efficiency. The average solar flux is -min 800wm-2 .water enters the collector at 30°C and flows at 2 litre min -1. what is the waters outlets temperature.


1
Expert's answer
2021-05-10T08:27:41-0400

"\eta \Phi=\frac{Q}{60St},"

"Q=c\rho v t(t_2^°-t_1^°),"

"60\eta \Phi S=c\rho v(t_2^°-t_1^°),"

"t_2^°=t_1^°+\frac{60\eta \Phi S}{c\rho v},"

"t_2^°=30+\frac{60\cdot 0.7\cdot 800\cdot1.2\cdot 2.4}{4200\cdot 1000\cdot 2}=30.012°C."